# Competing Mechanisms at Vibrated Interfaces of Density-Stratified Fluids

### v1

@@ Abstract
We focus on the case where both the RT and Faraday instabilities coexist, achieved by considering a negative gravity ($g^{*}=-1$). Figure 2 shows the growth rates of the displacement of the interface for various combinations of parameters. The lighter fluid properties are chosen to approximate three types of interfaces: inviscid liquid–liquid ($\mathrm{At}=0.3$, $\eta=0.1$), liquid–liquid ($\mathrm{At}=0.3$, $\eta=1$), and gas–liquid ($\mathrm{At}=0.9$, $\eta=0.1$). In the absence of vibration, $a^{*}=0$, classical RT instability manifests as a low-wavenumber peak. As the oscillation amplitude increases, the amplitude of the positive peak diminishes toward zero, while its position shifts toward the origin. Within a finite horizontal domain, this leads to the truncation of the RTs wave with $k^{*}\to 0$. This behavior demonstrates the theoretical principle of dynamic stabilization of RT waves under vibration [27, 28]. Given its continuous evolution with increasing vibration, we refer to the instability mechanism of the low-wavenumber peak as RT-type. The Floquet exponent associated with the RT peak is purely real and can thus be generalized to the cases of harmonic response. The Faraday instability emerges as the vibration amplitude increases, marked by a distinct peak at higher wavenumbers that shifts downward. The increasing prominence of the Faraday peak does not resonate with the RT peak. Instead, the RT peak is suppressed and pushed towards the zero-wavenumber limit, further highlighting the competitive interplay between these two phenomena. For all oscillating frequencies considered, the boundary layer thickness, $\delta\equiv\sqrt{2\nu_{d}/\omega}$, remains much lower than the fluid depth, with a maximum ratio of $\delta/h\leq 0.027$. In this regime, the onset of the Faraday instability is subharmonic [24]. While an increase in the viscosity ratio, $\eta$, has little effect on the growth rate profiles, increasing the density ratio, $\mathrm{At}$, induces local peaks at higher harmonics. Unstable higher-order harmonics occur at the lowest oscillation frequency ($\omega^{*}/2\pi=1.82$) in fig. 2 (d, g), complementing the dominant Faraday subharmonic mode. As expected, increasing the oscillation magnitude will eventually cause the Faraday peak to surpass the RT peak. This transition between the two fundamental hydrodynamic instabilities leads to a mixed region where both instabilities coexist, potentially altering the dynamics of the interface.
Figure 3: Multi-modal instability analysis for $(\mathrm{At},\eta)=(0.9,0.1)$: (a) dominant growth rate in the oscillating frequency-amplitude phase space, with the margins for the Faraday stability and the RT-Faraday transition highlighted; (b,c) normalized Faraday and RT wave evolutions with growth rates (insets) for cases marked by green crosses in (a), under single-mode or Perlin noise excitations. The wavenumber–time diagram (d) and not-to-scale interface profiles (e) show the nonlinear damping of Faraday waves for the Perlin noise case in (c).
Figure 3 (a) shows the dominant growth rate within the oscillating frequency-amplitude phase space. The transition of the leading instability mechanism from RT to Faraday can be observed in three stages. At low oscillation amplitudes, the RT instability is the dominant phenomenon. For larger oscillation amplitudes, the Faraday instability approaches its unstable boundary, yet the RT mechanism is dominant. Finally, at sufficiently high oscillation amplitudes, the growth rate of the Faraday instability surpasses that of the RT instability. For a given oscillation frequency, the thresholds of these two stages differ by about $2a^{*}$.
@@
To quantitatively compare theoretical predictions and numerical simulations, we perturb the initial interface with three profiles: single-mode RT and Faraday sine waves and multiscale Perlin noise perturbations [41]. The temporal dynamics of the normalized RT and Faraday modes are shown in fig. 3 (b, c), corresponding to the two cases marked in fig. 3 (a). While classical linear stability analysis of the static RT problem predicts purely exponential modal growth, the presence of vibrations induces an oscillatory pattern around this baseline growth. Owing to its subharmonic nature, the Faraday displacement becomes negative at $t^{*}=T^{*}$. The interface rebounds to a positive peak that either intersects the RT displacement (panel b) or exceeds it (panel c) at $t^{*}=2T^{*}$, where $T^{*}\equiv 2\pi/\omega^{*}$. For perturbations with a single sine wave, the modal growth of the most unstable Floquet exponent, $\mathrm{e}^{\gamma^{*}_{U}t^{*}}$, exhibits favorable agreement with the Fourier components obtained from simulations at integer multiples of the oscillation period for both instabilities. This alignment confirms that the interface growth follows Floquet theory. By incorporating the modal growth with the additional periodic terms, the total growth defined in Eq. 2 accurately predicts the interface dynamics during the first oscillation period, even with a broadband initial wavenumber spectrum. Beyond this period, the theory remains in qualitative agreement with simulations, although minor deviations arise due to nonlinear effects. Simulations initialized with Perlin noise exhibit suppressed growth rates and reduced amplitudes of Faraday waves compared to theoretical predictions. In the absence of competing wavelengths, these waves would exhibit Floquet-type growth and ultimately induce interface breakup, suggesting that the observed suppression is primarily attributable to nonlinear damping effects. When the Faraday growth rate exceeds that of RT, the theory remains qualitatively predictive of the system’s dynamics, despite the presence of multiscale waves. In contrast, the growth of RT waves is generally unaffected.
@@
To illustrate transient interface dynamics across wavenumbers in a realistic context, fig. 3(d) presents the wavenumber–time diagram for the Perlin noise case shown in fig. 3(c). This case exemplifies the nonlinear damping of Faraday waves despite a dominant linear growth rate. The coexistence of RT and Faraday waves is clearly observed within the given time window. Wavenumbers near the Faraday peak exhibit periodic phase changes due to their dominance of the subharmonic nature, whereas those near the RT peak display steady growth. The insets of fig. 3 (d) show the resulting interface profiles and highlight the competition between steadily amplifying long-wavelength RT waves and oscillatory short-wavelength Faraday waves. Initially ($t\lesssim 2.5T^{*}$), the interface profile is dominated by Faraday waves due to their higher initial growth rate. As time progresses, long-wavelength RT waves amplify and prevail, though residual small-scale Faraday oscillations persist. This transition implies that the growing low-wavenumber RT components nonlinearly dampen the high-wavenumber subharmonic Faraday responses, despite the inherent instability of these latter responses. As shown in fig. 3 (b, c), the Faraday waves would otherwise grow unimpeded toward interface breakup in the absence of RT competition. The dominance of RT waves suppresses this pathway, redirecting the system toward RT-driven dynamics. This mechanism has been verified by exciting the interface using only the linear superposition of Faraday and RT waves (not shown). Together, these results demonstrate the nonlinear transient dynamics that govern fluid interfaces under the combined effects of density and vibration.
IVConclusion
@@
This letter presents a theoretical and computational investigation into the previously unexplored coexistence and competition of two fundamental hydrodynamic instabilities: the pressure gradient-driven Rayleigh–Taylor (RT) instability and the resonance-induced Faraday instability. Through linear Floquet stability analysis and 2D numerical simulations, we demonstrate that interface growth under combined density and vibration effects exhibits a Floquet-type behavior: total displacement decomposes into modal growth of the most unstable Floquet mode and harmonic oscillations from periodic terms. Our key finding is the discovery of a novel multi-modal instability region, where the RT and Faraday instabilities coexist and compete in non-trivial ways. This region emerges as the oscillation magnitude increases, marking a gradual transition in the dominant instability mechanism from RT to Faraday. Within this region, we observed a unique competitive interaction: the Faraday instability consistently suppresses the RT instability toward the zero-wavenumber limit, while the residual RT mechanism, in turn, attenuates the Faraday responses via a nonlinear damping mechanism, even when the initial growth rate of the Faraday instability exceeds that of RT. Furthermore, the coexistence of these mechanisms generates multiscale interfacial waves, culminating in interfacial breakup phenomena that are ubiquitous in both natural and engineered systems, yet remain uncharted in prior studies.
@@
The discovery of the coexistence and competition between RT and Faraday instabilities offers new insights into predicting and controlling fluid interfaces in vibration-prone systems. Our findings bridge fundamental hydrodynamics and engineering, leveraging instability interactions of the fluid to control the interface kinematics. Vibrations dynamically stabilize RT modes while exciting Faraday instabilities, and the inclusion of RT waves, in turn, nonlinearly dampens the unstable Faraday response. By tuning the oscillation, processes such as atomization or inertial confinement fusion can be optimized to suppress both low- and high-wavenumber instabilities, thereby enhancing efficiency and preventing failure. These insights also open up new avenues for studying transient dynamics in both natural and engineered fluid systems, where external vibrations compete with intrinsic instabilities, ranging from geophysical flows to the injection of bubbles.

### v2

@@ III Numerical method
Here, we use the six-equation diffuse interface model [42] to describe the multiphase system, which we augmented with a surface tension formulation [43]. The governing equations have the form
$\partialderivative{{\bf\it q}}{t}+\gradient\cdot{\bf\it F}({\bf\it q})+{\bf\it h}({\bf\it q})\gradient\cdot{\bf\it u}={\bf\it s}({\bf\it q}),$ (23)
for which
[…]
For each fluid $j$, $\alpha_{j}$ is the volume fraction, $e_{j}$ is the internal energy, and
${\bf\it T}_{j}\equiv\mu_{j}\left[\gradient{\bf\it u}+(\gradient{\bf\it u})^{\mathsf{T}}-\frac{2}{3}\left(\gradient\dotproduct{\bf\it u}\right){\bf\it I}\right]$ (25)
is the viscous stress tensor. The equations are closed by the usual set of mixture rules
$\sum_{j}\alpha_{j}=1,\quad\rho=\sum_{j}\rho_{j}\alpha_{j},\quad\mu=\sum_{j}\mu_{j}\alpha_{j},\quad\text{and}\quad{\bf\it T}=\sum_{j}{\bf\it T}_{j}\alpha_{j}.$ (26)
The total mixture energy $E$ is
[…]
where $Y_{j}=\alpha_{j}\rho_{j}/\rho$ are the mass fractions of each phase. Surface tension and body forces are applied only to the total mixture energy. The capillary stress tensor $\bm{\Omega}$ is given by
$\bm{\Omega}=-\sigma\left(\lVert\nabla\alpha_{1}\rVert\bm{I}-\frac{\nabla\alpha_{1}\otimes\nabla\alpha_{1}}{\lVert\nabla\alpha_{1}\rVert}\right),$ (28)
and $\epsilon_{0}=\sigma\lVert\nabla\alpha\rVert$ is the capillary energy. The system is solved via a high-order accurate finite-volume method and Runge–Kutta temporal discretization. The numerical method closely follows that of Coralic and Colonius [44] and is described in detail in Bryngelson et al. [39] and Wilfong et al. [41]. Closure is provided by the stiffened gas equation of state
$e_{j}=\frac{p_{j}+\gamma_{j}\pi_{\infty,j}}{\left(\gamma_{j}-1\right)\rho_{j}}.$ (29)
The fluid parameters for ratio of specific heats, $\gamma_{j}$, and liquid stiffness, $\pi_{\infty,j}$, enable faithful modeling of liquids and gasses [45].
Numerical simulations   The 2D computational domain spans four wavelengths of interest in width, with each fluid layer having a depth of $h=$0.013\text{\,}\mathrm{m}$\$. A uniform Cartesian mesh resolves the shortest wavelength with 192 grid cells, which is sufficient to accurately capture the interface between the two fluids. This resolution has been validated through a grid convergence study conducted with up to 288 grid cells. Periodic and no-slip boundary conditions are applied at the horizontal and vertical boundaries, respectively. The interface is initially perturbed by a height that is $1\%$ of the wavelength, and hydrodynamic equilibrium is achieved at the interface with a pressure of $100\text{\,}\mathrm{kPa}\text{/}$. For demonstration purposes, we consider polydimethylsiloxane (PDMS) as the more dense fluid, which has been used in experimental studies of vibrated fluids [46]. PDMS is characterized as a Newtonian fluid with density $\rho_{d}=$950\text{\,}\mathrm{kg}\text{/}{\mathrm{m}}^{3}$\$, kinematic viscosity of $\nu_{d}=$2e{-5}\text{\,}{\mathrm{m}}^{2}\text{/}\mathrm{s}$\$, and surface tension coefficient of $\sigma_{d}=$2.06e{-2}\text{\,}\mathrm{N}\text{/}\mathrm{m}$\$. The stiffened gas equation of state parameters $(\gamma,\pi_{\infty})=(3.49,2.78\times 10^{8})$ for PDMS are selected to match the speed of sound $1011\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$ [47] and the internal energy of water at standard temperature and pressure. The gravitational acceleration is $g=$9.81\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}^{2}$\$, yielding the dimensionless quantities $C=0.0043$, and $\mathrm{Bd}=76.456$, which quantify the ratios of viscous and surface tension forces to gravitational forces, respectively. The properties of the lighter fluid are determined by the density and viscosity ratios, $\mathrm{At}$ and $\eta$. Table 2 summarizes the configurations studied. The equation of state parameters for the lighter fluid are $(\gamma,\pi_{\infty})=(1.4,0)$ for the $\mathrm{At}=1$ case and $(\gamma,\pi_{\infty})=(3.48,1.46\times 10^{7})$ for the $\mathrm{At}=0.9$ case. A startup period lasting a quarter of an oscillation cycle initiates fluid motion, ensuring consistency with the non-zero velocity fields predicted by linear theory. Phase interfaces are identified by isolating the $\alpha=0.5$ isosurface of the volume fraction field.
@@
Figure 5 (a) 13.00 16 1-1 33.9 (0.9,0.1) 2.89 1.21 (H)
13.00 16 1-1 33.9 (0.9,0.1) 29.1 1.21 (S)
Figure 5 (b) 13.00 18 1-1 27.4 (0.9,0.1) 2.04 0.96 (H)
13.00 18 1-1 27.4 (0.9,0.1) 24.5 3.45 (S)
IV Results
We verify the linear theory predictions for the RT-stable case, where $g^{*}=1$. In this scenario, the interface is modeled as an ideal free surface, with the lighter fluid assumed to have zero density and viscosity. Consequently, only Faraday-type instabilities can be triggered via vertical vibration. Figure 2 (a) shows the growth rates for four oscillation amplitude and frequency combinations, each with a positive maximum growth rate at wavenumber $k^{*}/2\pi=5$. This behavior indicates the onset of Faraday instability, leading to an unstable interface. For all oscillation frequencies considered, the boundary layer thickness, $\delta\equiv\sqrt{2\nu_{d}/\omega}$, is sufficiently smaller than the fluid depth ($\delta/h\leq 0.019$) and so the onset of the Faraday instability is subharmonic [25]. The temporal evolution of the resulting Faraday waves over three oscillation periods is shown in fig. 2 (b–e). Initial transient dynamics of the Faraday waves are faithfully represented by the exponential modal growth of the most unstable Floquet exponent and its associated periodic behavior, with less than 1% error compared to numerical simulations. This agreement suggests that the nonmodal growth of Faraday waves is effectively represented. These Faraday waves grow subharmonically and eventually lead to interface breakup.
Refer to captionParameters are provided in table 2
Figure 2: Analysis of cases with the most unstable wavenumber $k^{*}/2\pi=5$ for $g^{*}=1$: (a) Growth rates; (b–e) Temporal evolution of the most unstable waves, with a magnified view showing the first period.
@@ IV Results
We focus on the case where both the RT and Faraday instabilities coexist, specifically under a gravitational configuration directing gravity from the denser fluid toward the lighter one ($g^{*}=-1$). Figure 3 shows the growth rates of the displacement of the interface for various combinations of parameters. The lighter fluid properties are chosen to approximate three types of interfaces: inviscid liquid–liquid ($\mathrm{At}=0.3$, $\eta=0.1$), liquid–liquid ($\mathrm{At}=0.3$, $\eta=1$), and gas–liquid ($\mathrm{At}=0.9$, $\eta=0.1$). A classical RT instability manifests as a low-wavenumber peak in the zero-vibration case $a^{*}=0$. For larger oscillation amplitudes, the amplitude of the positive peak diminishes toward zero and its position shifts toward the coordinate origin. This suppression does not lead to a complete stabilization of the RT instability in infinite domains, as the RT growth rate $\gamma_{U}^{*}$ remains positive, albeit reduced, under vibration. In contrast, complete stabilization can be achieved in a finite horizontal domain, where the truncation of long-wavelength RT modes with $k^{*}\to 0$ effectively eliminates their growth. This behavior demonstrates the theoretical principle of dynamic stabilization of RT waves under vibration within a confined container [31, 32]. Given its continuous evolution with increasing vibration, we refer to the instability mechanism of the low-wavenumber peak as RT-type. The Floquet exponent associated with the RT peak is purely real and can thus be generalized to the cases of harmonic response. The Faraday instability emerges as the vibration amplitude increases, marked by a distinct peak at higher wavenumbers that shifts downward. The increasing prominence of the Faraday peak does not resonate with the RT peak. Instead, the RT peak is suppressed and pushed towards the zero-wavenumber limit, further highlighting the competitive interplay between these two phenomena. While an increase in the viscosity ratio, $\eta$, has little effect on the growth rate profiles, increasing the density difference, $\mathrm{At}$, induces local peaks at higher harmonics. Unstable higher-order harmonics occur at the lowest oscillation frequency ($\omega^{*}/2\pi=1.82$) in fig. 3 (d, g), complementing the dominant Faraday subharmonic mode. As expected, increasing the oscillation amplitude will eventually cause the Faraday peak to surpass the RT peak. This transition between the two fundamental hydrodynamic instabilities leads to a mixed region where both instabilities coexist, potentially altering the dynamics of the interface.
Figure 4: Dominant growth rate across the oscillating frequency-amplitude phase space: (a) $(\mathrm{At},\eta)=(0.9,0.1)$; (b) $(\mathrm{At},\eta)=(0.3,0.1)$. The margins for the Faraday instability and the RT-Faraday transition are highlighted.
To have a comprehensive understanding of the multi-modal instability region and the associated transition mechanisms, fig. 4 shows the dominant growth rate within the oscillating frequency-amplitude phase space for two different density ratios. For both cases, the transition of the leading instability mechanism from RT to Faraday can be observed in three stages. At low oscillation amplitudes, the RT instability is the dominant phenomenon. The Faraday instability approaches its unstable boundary for larger oscillation amplitudes, yet the RT mechanism is dominant. Finally, at sufficiently high oscillation amplitudes, the growth rate of the Faraday instability surpasses that of the RT instability. For a given oscillation frequency, the thresholds of these two stages differ by about $2g^{*}$ for $\mathrm{At=0.9}$. When the density difference between the two fluids is reduced ($\mathrm{At}=0.3$), the RT instability dominates most of the phase space under consideration, albeit with a reduced growth rate. The onset of Faraday instability requires higher oscillation amplitudes at the same oscillation frequency. The threshold separating RT-dominated and Faraday-dominated regimes decreases to approximately $0.5g^{*}$. These trends are consistent with what one expects; both instabilities weaken in the limit of vanishing density contrast. In the following analysis, we focus on the case of $\mathrm{At}=0.9$, where both instabilities are more strongly amplified compared to cases with lower density contrast.
Figure 5: Normalized evolution of Faraday and RT waves with growth rates (insets) under single-mode or Perlin noise excitations: (a) $a^{*}=16,\,\omega^{*}=33.9$; (b) $a^{*}=18,\,\omega^{*}=27.4$. Both are marked in fig. 4 (a).
@@
To quantitatively compare theoretical predictions and numerical simulations, we perturb the initial interface with three profiles: single-mode RT and Faraday sine waves and multiscale Perlin noise perturbations [48]. The temporal dynamics of the normalized RT and Faraday modes are shown in fig. 5  (a, b), corresponding to the two cases marked in fig. 4 (a). While classical linear stability analysis of the static RT problem predicts purely exponential modal growth, vibrations induce an oscillatory pattern around this baseline growth. Owing to its subharmonic nature, the Faraday displacement becomes negative at $t^{*}=T^{*}$. The interface rebounds to a positive peak that either intersects the RT displacement (panel a) or exceeds it (panel b) at $t^{*}=2T^{*}$, where $T^{*}\equiv 2\pi/\omega^{*}$. For perturbations with a single sine wave, the modal growth of the most unstable Floquet exponent, $\mathrm{e}^{\gamma^{*}_{U}t^{*}}$, exhibits favorable agreement with the Fourier components obtained from simulations at integer multiples of the oscillation period for both instabilities. This alignment confirms that the interface growth follows Floquet theory. By incorporating the modal growth with the additional periodic terms, the total growth defined in Eq. 11 accurately predicts the interfacial dynamics during the first oscillation period. Beyond this period, the theory remains in qualitative agreement with simulations, although minor deviations arise due to nonlinear effects. Simulations initialized with Perlin noise show suppressed growth rates and reduced amplitudes of Faraday waves compared to theoretical predictions. The broadband initial wavenumber spectrum induces a harmonic-like response in the Faraday waves when their growth rate matches that of the RT instability. Without competing wavelengths, these waves would exhibit subharmonic Floquet-type growth and ultimately induce interface breakup, suggesting that the observed suppression is primarily attributable to nonlinear damping effects. When the Faraday growth rate exceeds that of RT, the theory remains qualitatively predictive of the system’s dynamics, despite multiscale waves. In contrast, the growth of RT waves is generally unaffected by the presence of multiscale wave components.
Figure 6: The wavenumber–time diagram for the Perlin noise case in fig. 5 (b).
@@
To illustrate transient interface dynamics across wavenumbers in a realistic context, fig. 6 presents the wavenumber–time diagram for the Perlin noise case shown in fig. 5 (b). This case exemplifies the nonlinear damping of Faraday waves despite a dominant linear growth rate. The coexistence of RT and Faraday waves is observed within the given time window. Wavenumbers near the Faraday peak exhibit periodic phase changes due to their dominance of the subharmonic nature, whereas those near the RT peak display steady growth. Initially ($t\lesssim 2.5T^{*}$), the interface profile is dominated by Faraday waves due to their higher initial growth rate. Long-wavelength RT waves amplify and dominate as time progresses, though residual small-scale Faraday oscillations persist. This transition implies that the steadily growing RT components nonlinearly dampen the oscillatory subharmonic Faraday responses, despite the inherent instability of these latter responses. The resulting interface profiles, shown later in fig. 7, provide a direct visualization of the competition between long-wavelength RT waves and short-wavelength Faraday waves. Figure 5 (a, b) shows that the Faraday waves would otherwise grow unimpeded toward interface breakup without RT competition. The dominance of RT waves suppresses this pathway, redirecting the system toward RT-driven dynamics. Together, these results demonstrate the nonlinear transient dynamics that govern fluid interfaces under the combined effects of density and vibration.
Figure 7: Temporal evolution of the vertical momentum field and interface profile.
To better understand the nonlinear damping of Faraday waves that cannot be represented via linear theory, we examine the evolution of the entire flow field, rather than the interface alone. Figure 7 shows the vertical momentum field, $\rho^{*}w^{*}$, a conservative quantity, at various time instants. A reference case is also considered for direct comparison, in which a linear superposition of Faraday and RT modes with equal amplitude perturbs the initial interface. The velocity fields exhibit wavenumber patterns consistent with those observed at the interface. Faraday-type wave packets remain localized near the interface, while RT-type wave packets penetrate deeper, occupying the entire lower half of the domain. These spatial structures reflect their underlying physical mechanisms: the former arises from interfacial resonance, while buoyancy-induced pressure gradients drive the latter. Faraday waves are readily developed in the absence of competing length scales. Their nonlinear damping does not require broadband perturbations such as Perlin noise. Including a single additional RT mode is sufficient to suppress the Faraday response. As discussed in the context of fig. 3, increasing the vibration amplitude to amplify the Faraday instability will consistently suppress the RT instability toward the zero-wavenumber limit and reduce its growth rate. Numerical simulations further reveal that the residual, weakened RT dynamics concurrently attenuate the Faraday response through a nonlinear damping mechanism. These results demonstrate a bidirectional competitive interaction between the two fundamental instabilities, each modulating the other through linear and nonlinear processes.
V Conclusions
@@
This manuscript presents a theoretical and computational investigation into the previously unexplored coexistence and competition of two fundamental hydrodynamic instabilities: the pressure gradient-driven Rayleigh–Taylor (RT) instability and the resonance-induced Faraday instability. Through linear Floquet stability analysis and 2D numerical simulations, we demonstrate that interface growth under combined density and vibration effects exhibits a Floquet-type behavior: total displacement decomposes into modal growth of the most unstable Floquet mode and harmonic oscillations from periodic terms. In an unbounded horizontal domain, we show that vibration can mitigate the growth rate of RT instabilities. This result aligns with the dynamical stabilization of RT instabilities in confined containers, where physical boundaries restrict long-wavelength modes. We also reveal a previously unseen multi-modal instability region, where the RT and Faraday instabilities coexist and compete in non-trivial and unintuitive ways. This region emerges as the oscillation amplitude increases, marking a gradual transition in the dominant instability mechanism from RT to Faraday. Within this region, we observed a unique bidirectional competitive interaction. The growing Faraday mechanism, amplified by increasing vibration amplitude, consistently suppresses the RT instability toward the zero-wavenumber limit. In turn, the residual RT mechanism attenuates the Faraday responses via a nonlinear damping mechanism, even when the initial growth rate of the Faraday instability exceeds that of RT. Furthermore, the coexistence of these mechanisms generates multiscale interfacial waves, culminating in interfacial breakup phenomena that are ubiquitous in both natural and engineered systems, yet remain uncharted in prior studies.
Although linear stability analysis predicts a multi-modal instability region where RT and Faraday instabilities should coexist, numerical simulations reveal a critical nuance. While each instability develops independently at its theoretically predicted growth rate, sustained coexistence is difficult to achieve due to the nonlinear damping of Faraday waves by RT-induced flows. This bidirectional competition offers a plausible explanation for the absence of simultaneous RT and Faraday instabilities in most natural settings. Nonlinear approaches, such as weakly nonlinear stability analysis, offer a promising way to account for these cross-mode interactions. Extending the analysis and simulations to 3D, although more computationally demanding, enables the capture of more complex interfacial instability patterns, such as squares and hexagons. These structures provide a more faithful representation of real-world scenarios. Nevertheless, the current 2D approach effectively captures the competition between the instability mechanisms at different scales and remains valid under geometric symmetries that permit dimensionality reduction, such as azimuthal decomposition in cylindrical containers.
@@
The discovery of coexistence and competition between RT and Faraday instabilities offers new insights into predicting and controlling fluid interfaces in vibration-prone systems. Our findings bridge fundamental hydrodynamics and engineering, leveraging instability interactions of the fluid to control the interface kinematics. Vibrations dynamically stabilize RT modes while exciting Faraday instabilities, and the inclusion of RT waves, in turn, nonlinearly dampens the unstable Faraday response. By tuning the oscillation, processes like atomization or inertial confinement fusion can be optimized to suppress both low- and high-wavenumber instabilities, thereby enhancing efficiency and preventing failure. These insights reveal new avenues for studying transient dynamics in natural and engineering flow systems, where external vibrations compete with intrinsic instabilities, ranging from geophysical flows to the injection of bubbles.